OpenStudy (anonymous):

Whats the correct slope of the line that passes through the points (−4, 8) and (−3, −6). please

OpenStudy (anonymous):

slope = rise/run = (y2 -y1)/(x2 -x1) =(-6-8) / (-3 -(-4)) =-14 / 1 = -14

OpenStudy (amistre64):

(-4, 8) -(-3,-6) --------- (-1,14) ; slope = y/x = -14 ; winks right ;)